Debye temperature of diamond is 2000k determine the mean velocity of diamond if the interatomic spacing is 1.45A. estimate the mode of a lattice vibration

1
Expert's answer
2022-01-18T09:36:20-0500

"h\nu=kT,\implies \nu=\frac{kT}{h},"

"\nu=416\cdot 10^{11}~Hz,"

"v=\nu l=6~\frac{km}s."